If a = 3i_4j & b=i- 4j calculate a bar+b bar and mod of a bar+ b bar

Grade:11

1 Answers

Vikas TU
7 years ago
a bar is the conjugate of 3i – 4j
thus abar = -3i + 4j
and bbar = -i + 4j
modulus would be first add them,
= |-4i +8j|
= root(16 + 64)
= root(80)